OpenStudy (anonymous):

d/dx (f(2x^2)) = 7x^4 Find f'(x)

OpenStudy (anonymous):

\[\frac{ d }{ dx } f(2x^2))=7x^4\]

zepdrix (zepdrix):

Hmmmmmmmmmmmmm...

zepdrix (zepdrix):

This is alongs the lines that I'm thinking...\[\Large f'(2x^2)\quad=\quad 7x^4\quad=\quad C(2x^2)^2\]

OpenStudy (anonymous):

Wait thats the wrong problem.......

zepdrix (zepdrix):

oh lol

OpenStudy (anonymous):

Ill post new 1 its just a few numbers changed

OpenStudy (anonymous):

\[d/dx (f(2x^4))=7x^3\]

OpenStudy (anonymous):

first step i got was \[f'(2x^4)=7x^3\]

OpenStudy (anonymous):

f' (2x^4)*8x^3 = 7x^3

OpenStudy (anonymous):

\[f'(2x^4)=\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

Im there.... i need to find f'(x)

zepdrix (zepdrix):

So you chain ruled on the left? Hmm looks like you're on the right track!

zepdrix (zepdrix):

So you determined that the derivative function is `constant`. So regardless of what value you plug into the function, 2x^4, x, a bajillion, it will always come back as 7/8. \[\Large f'(2x^4)\quad=\quad\frac{7}{8}\]\[\Large f'(200x^9)\quad=\quad\frac{7}{8}\]\[\Large f'(x)\quad=\quad?\]

OpenStudy (anonymous):

still 7/8

zepdrix (zepdrix):

yessss.. i think that's what we're looking for at least +_+

OpenStudy (anonymous):

O.O that was so easy

OpenStudy (anonymous):

can you help me with this word problem?

zepdrix (zepdrix):

i can try :u

OpenStudy (anonymous):

A spherical balloon is inflated so that its volume is increasing at the rate of 2.7 ft^3/min. How rapidly is the diameter of the balloon increasing when the diameter is 1.8 feet?

OpenStudy (anonymous):

The diameter is increasing at____ft/min

zepdrix (zepdrix):